# Lightweight Transformer via Unrolling of Mixed Graph Algorithms for Traffic Forecast

### v1

@@ G.1 Sensitivity of $k$, $W$ in Graph Construction
8 6 45.37 18.15 20.59 6 8 12.44 5.3525 11.96
G.2 Sensitivity of Number of Blocks and Layers in the Unrolling Model
@@
G.3 Ablation Study for the Splitting-$\ell_{2}$-term Strategy
We also ran ablations of the splitting-term strategy by optimizing ${\mathbf{x}}^{\tau+1}$ using only (7). CG in solving (7) converges much slower than in solving (10)-(12), and the unrolled model can become numerically unstable in training, indicating that splitting the $\ell_{2}$ terms can stabilize the training process.
Appendix H Error Bar for Our Model in Table 1
@@
Dataset PeMS03 METR-LA PEMS-BAY
$2\sigma_{\text{RMSE}}$ 0.05 0.04 0.01
$2\sigma_{\text{MAE}}$ 0.02 0.02 0.01
$2\sigma_{\text{MAPE(\%)}}$ 0.03 0.04 0.10
Appendix I Reconstruction Error
We perform the spatial/temporal forecast by recovering the entire time-vertex signal, which is unusual compared to other deep models. Table 8 shows that the reconstruction error is much smaller than the predicted part, indicating that our model has extracted all spatial/temporal correlations from the observed part for prediction.
Table 8: Reconstruction RMSE for the observed signals in 60-minute forecast in all three datasets.
Dataset PEMS03 METR-LA PEMS-BAY
rec RMSE 1.44 0.49 0.39

### v2

@@ H.2 Sensitivity of $k$, $W$ in Graph Construction
8 6 45.37 18.15 20.59 6 8 12.44 5.35 11.96
H.3 Sensitivity of Number of Blocks and Layers in the Unrolling Model
@@
6 25 27.84 17.54 18.61 12.37 5.27 11.90
H.4 Ablation Study for the Splitting-$\ell_{2}$-term Strategy
We also ran ablations of the splitting-term strategy by optimizing ${\mathbf{x}}^{\tau+1}$ using only eq. 7. CG in solving eq. 7 converges much slower than in solving eq. 10-eq. 12, and the unrolled model can become numerically unstable in training, indicating that splitting the $\ell_{2}$ terms can stabilize the training process.
Appendix I Optimization Formulation and ADMM Algorithm in Ablation Studies
I.1 Optimization & Algorithm without DGTV Term
By removing the DGTV term $\|{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{1}$ in eq. 4, and introducing the auxiliary variables ${\mathbf{z}}_{u},{\mathbf{z}}_{d}$ as eq. 9, we rewrite the unconstrained version of the optimization by augmented Lagrangian method as
$\min_{{\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{d}}$ $\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+\mu_{d,2}{\mathbf{z}}_{d}^{\top}{\mathcal{L}}^{d}_{r}{\mathbf{z}}_{d}+({\bm{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$
$+({\bm{\gamma}}_{d}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{d})+\frac{\rho_{d}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{d}\|_{2}^{2}.$ (68)
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\bm{\gamma}}_{u}^{\tau}+{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{d}}{2}{\mathbf{z}}_{d}^{\tau},$ (69)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (70)
$\left(\mu_{d,2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{d}}{2}{\mathbf{I}}\right){\mathbf{z}}_{d}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{d}^{\tau}}{2}+\frac{\rho_{d}}{2}{\mathbf{x}}^{\tau+1}.$ (71)
The updating of ${\bm{\gamma}}_{u}^{\tau},{\bm{\gamma}}_{d}^{\tau}$ follows eq. 18.
I.2 Optimization & Algorithm without DGLR Term
By removing the DGLR term ${\mathbf{x}}{\mathcal{L}}^{d}_{r}{\mathbf{x}}$ in eq. 4 and introduce only the auxiliary variables $\boldsymbol{\phi}$} in eq. 5 and ${\mathbf{z}}_{u}$ in eq. 9, we rewrite the unconstrained version of the optimization of ${\mathbf{x}}^{\tau+1}$ by augmented Lagrangian method as
$\min_{x,{\mathbf{z}}_{u}}\|{\mathbf{y}}-{\mathbf{H}}{\mathbf{x}}\|_{2}^{2}+\mu_{u}{\mathbf{z}}_{u}^{\top}{\mathbf{L}}^{u}{\mathbf{z}}_{u}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}})+\frac{\rho}{2}\|{\boldsymbol{\phi}}^{\tau}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}\|_{2}^{2}+({\bm{\gamma}}_{u}^{\tau})^{\top}({\mathbf{x}}-{\mathbf{z}}_{u})+\frac{\rho_{u}}{2}\|{\mathbf{x}}-{\mathbf{z}}_{u}\|_{2}^{2}$ (72)
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho}{2}{\mathcal{L}}^{d}_{r}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}$ $={\mathbf{H}}^{\top}{\mathbf{y}}+({\mathbf{L}}^{d}_{r})^{\top}(\frac{\gamma^{\tau}}{2}+\frac{\rho}{2}{\boldsymbol{\phi}}^{\tau})-\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}$ (73)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}$ $=\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1}$ (74)
The optimization of ${\boldsymbol{\phi}}^{\tau+1}$ is the same as eq. 15, and the solution is the same as eq. 16. The updating of ${\bm{\gamma}},{\bm{\gamma}}_{u}$ follows eq. 17 and eq. 18.
I.3 Algorithm Unrolling with an Undirected Temporal Graph
where $\mu_{u},\mu_{n}\in\mathbb{R}_{+}$ are weight parameters for the two GLR priors. The opitimization is similar to that in Appendix I.1. By a similar approach which introduces auxiliary variables ${\mathbf{z}}_{u}={\mathbf{x}},{\mathbf{z}}_{n}={\mathbf{x}}$, we optimize ${\mathbf{x}},{\mathbf{z}}_{u},{\mathbf{z}}_{n}$ as follows:
$\left({\mathbf{H}}^{\top}{\mathbf{H}}+\frac{\rho_{u}+\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{x}}^{\tau+1}=$ ${\mathbf{H}}^{\top}{\mathbf{y}}-\frac{{\bm{\gamma}}_{u}^{\tau}+{\bm{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{z}}_{u}^{\tau}+\frac{\rho_{n}}{2}{\mathbf{z}}_{n}^{\tau},$ (76)
$\left(\mu_{u}{\mathbf{L}}^{u}+\frac{\rho_{u}}{2}{\mathbf{I}}\right){\mathbf{z}}_{u}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{u}^{\tau}}{2}+\frac{\rho_{u}}{2}{\mathbf{x}}^{\tau+1},$ (77)
$\left(\mu_{n}{\mathbf{L}}^{n}+\frac{\rho_{n}}{2}{\mathbf{I}}\right){\mathbf{z}}_{n}^{\tau+1}=$ $\frac{{\bm{\gamma}}_{n}^{\tau}}{2}+\frac{\rho_{n}}{2}{\mathbf{x}}^{\tau+1}.$ (78)
The updating of ${\bm{\gamma}}_{u}^{\tau},{\bm{\gamma}}_{n}^{\tau}$ is also similar to eq. 18 as follows:
${\bm{\gamma}}_{u}^{\tau+1}={\bm{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),\quad{\bm{\gamma}}_{n}^{\tau+1}={\bm{\gamma}}_{n}^{\tau}+\rho_{n}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{n}^{\tau+1}).$ (79)
The undirected temporal graph ${\mathcal{G}}^{n}$ is learned by an UGL described in Section 4.3, and is initialized similar to the spatial UGL described in Appendix F.2. The initialization of $\mu_{n},\rho_{n}$ and the CGD parameters in solving eq. 78 follows the setup in Appendix F.1.
Appendix J Error Bar for Our Model in Table 2
@@ I.3 Algorithm Unrolling with an Undirected Temporal Graph
Dataset PEMS03 METR-LA
$2\sigma_{\text{RMSE}}$ 0.05 0.04
$2\sigma_{\text{MAE}}$ 0.02 0.02
$2\sigma_{\text{MAPE(\%)}}$ 0.03 0.04